OpenStudy (anonymous):

The sides of a quadrilateral are 3,4,5 and 6. Find the length of the shortest side of a similar quadrilateral whose area is 9 times as great. 9 13.5 27

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@driftracer305

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Zarkon

ganeshie8 (ganeshie8):

if area is \(\large k\) times greater, then the side will be \(\large \sqrt{k}\) times greater

OpenStudy (anonymous):

So the shortest side that we know is 3 so 3k

OpenStudy (anonymous):

So 27

OpenStudy (anonymous):

area goes with the square if the area is 9 times as great then each side is 3 times as large

OpenStudy (anonymous):

Oh ok thanks for taking the time to help me

OpenStudy (anonymous):

a simple example would be a square of side say 5 the area would be 25 then \(9\times 25= 225\) and if a square had an area 225 then the side would be length \(3\times 5=15\)

OpenStudy (anonymous):

hope it is clear yw

OpenStudy (anonymous):

Yeah that helped, do you mind helping me with one more

OpenStudy (anonymous):

i kind of suck at geometry but i could try

OpenStudy (anonymous):

if i can't help i will just say so

OpenStudy (anonymous):

Ok thanks. A circle has a radius of 6 in. The inscribed equilateral triangle will have an area of?

OpenStudy (anonymous):

i think it is \[\frac{\sqrt3}{4}r^2\]right?

OpenStudy (anonymous):



OpenStudy (anonymous):

unless you need a proof that should work

OpenStudy (anonymous):

Hmmidk for sure I'll try that out

OpenStudy (anonymous):

wait hold the phone that is wrong it is three times that

OpenStudy (anonymous):

I just wants the answer that is two digits times the square root of one for the answer

OpenStudy (anonymous):

It*

OpenStudy (anonymous):

damn typo \[\frac{\sqrt3}{4}\times 6^2\times 3\]

OpenStudy (anonymous):

Ahh ok thanks, too many formula to memorize and I'm having to teach this to myself

OpenStudy (anonymous):

better know as \(27\sqrt3\)

OpenStudy (anonymous):

i could explain it but not without trig

OpenStudy (anonymous):

Ok thanks for the help, I appreciate it

OpenStudy (anonymous):

yw